[Brahms-l] Shift report 20040211 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Wed Feb 11 2004 - 23:43:38 EST
------------------------------------------------------------
Date: Wed Feb 11 23:43:38 2004, Shift: 16:00-24:00
Supervisor: BHS, DO, with: A mouse in the ceiling
Log book#15 pages 53 - 53

Shift Summary:
------------------------------------------------------------
1: Beam exp. until 18.30, then beam 19.00-23.30. Low BG rates, nice store.

2: Good data runs:

Run       FS      Trig2   MRS       Trig3
10273 -   4,1/1A  110k    45,1/1B   100k
10275 -   4,1/2A  130k     "        140k
10276 -   4,1/2A   90k     "         90k
10277 -   4,1/2A   40k     "         40k

No DC trips :-)

3: Problems
1) D1 tripped after approx. 1 hour due to overheating, had to be reset by MagnaDad. To avoid this problem we went to 1/2 field, and MagWatch will try to fix this tomorrow so we can go back to full field. 
2) Found a strange hole in the Silicon Summary 2 plot (channel vs. ADC) - see e-log.

That's about it - RHIC is being prepared for another fill.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Feb 11 23:44:04 2004

This archive was generated by hypermail 2.1.8 : Wed Feb 11 2004 - 23:44:28 EST